# Mitigating Multimodal Hallucinations via Gradient-based Self-Reflection

### v3

@@ 1 Introduction
Recent advances in Multimodal Large Language Models (MLLMs) show strong ability to produce coherent and context-aware content across a wide range of domains [bai2023qwen, dai2023instructblipgeneralpurposevisionlanguagemodels, chen2024internvl, liu2024llava, ye2024mplug]. Despite their impressive advancements, these models remain prone to hallucination, wherein the generated text is not faithfully grounded in the visual modality [rohrbach2018object, li2023evaluating]. This limitation poses a critical barrier to establishing trust in the outputs of MLLMs.
@@
The hallucinations observed in MLLMs can be largely attributed to two fundamental biases [kang2023impact, li2023evaluating, kim2024discovering]. Text-visual bias refers to the excessive reliance on textual information—such as the input prompt and previously generated outputs—while neglecting the visual modality during generation. This bias becomes particularly pronounced in longer sequences, where MLLMs tend to depend more heavily on prior text and increasingly disregard visual cues [zhou2023analyzing, favero2024multi]. Co-occurrence bias arises from spurious statistical correlations embedded in the training data, which lead models to erroneously predict the presence of non-existent objects based on their frequent co-occurrence with observed objects in the visual inputs [li2023evaluating]. This bias is particularly challenging to mitigate, and existing approaches largely rely on statistical priors rather than offering statistically agnostic solutions [kang2023impact, zhou2023solution].
@@
Existing efforts to mitigate hallucinations in MLLMs can be broadly categorized into inference-based methods, which operate at the decoding stage [chen2024halc, favero2024multi, leng2024mitigating, park2024convis, woo2024don], and training-based approaches, which intervene during model optimization [ben2023mocha, chen2023mitigating, kang2023impact, sun2023aligning, jiang2024hallucination, yue2024less]. Inference-based approaches are valued for their cost-effectiveness, as they avoid the need for additional data collection, data bias examination, or extensive model retraining. However, these methods offer limited insight into the severity of underlying biases, leaving the root causes of hallucination insufficiently understood. In addition, some inference-based methods rely on auxiliary models—such as segmentation networks  [chen2024halc], detection systems  [kan2024catchcomplementaryadaptivetokenlevel] , or even additional MLLMs [radford2021learning, deng2024seeing, xing2024efuf]—which undermine their purported cost-effectiveness.
@@
Another limitation of existing methods lies in their lack of granularity when adjusting the underlying biases in MLLMs. Most approaches rely on heuristically tuned priors, which vary across datasets and fail to generalize reliably [leng2024mitigating, zhao2024mitigating]. Moreover, they apply uniform weighting across all visual features, offering no mechanism to selectively adjust bias at the level of individual features [zhang2024debiasing, manevich2024mitigatinghallucinationslargevisionlanguage]. This coarse treatment limits their effectiveness in mitigating co-occurrence bias, which arises from spurious statistical correlations between objects that are often represented by distinct visual features.
@@
  • Extensive experiments demonstrate that GACD mitigates hallucinations and enhances accuracy while maintaining a favorable balance between accuracy and informativeness. GACD achieves up to $8\%$ increase in overall score on AMBER [wang2023llm], an $8\%$ F1 boost on POPE [li2023evaluating], up to $45\%$ improvement in detailness and a $92\%$ accuracy gain on LLaVA-QA90 [liu2024visual].
@@ 2 Related Work
Hallucination and Bias. Hallucinations in LLMs often arise from biases in the training data [mckenna2023sources, huang2025survey], while in MLLMs, studies [tonmoy2024comprehensive, li2023evaluating, fu2024mmecomprehensiveevaluationbenchmark] show that hallucinations are closely linked to biases like text-visual and co-occurrence biases. Additionally, biases related to output position, which increase the risk of hallucination as output length grows, have been examined in [favero2024multi, zhou2023analyzing]. Existing methods [li2023evaluating, fu2024mmecomprehensiveevaluationbenchmark, kim2024discovering] typically report only overall statistics, lacking a mathematical, sample-wise bias measurement. This distinction is important, as biases can vary case by case. Our approach measures sample-dependent bias via token-level gradient sensitivities, revealing how pre-trained MLLM parameters embed these biases  [kim2019learning, guo2024bias], and enabling self-reflective hallucination mitigation.
Hallucination Mitigation. Training-related hallucination mitigation methods [chen2023minigpt, jiang2024hallucination, yue2024less, peng2025mitigating, zadeh2025lpoi] are expensive, requiring access to training data and specialized statistical analysis. Among them, LPOI [zadeh2025lpoi] also employs an object-aware framework, highlighting the effectiveness of modeling object-level information for mitigating hallucinations. Reinforcement‐learning approaches [xing2024efuf, deng2024seeing, zhai2024hallecontrolcontrollingobjecthallucination] rely on supplementary feedback, often from human annotators or auxiliary LLMs/MLLMs, and the latter may themselves hallucinate. By contrast, post-decoding techniques modify model logits at inference time without further training or external feedback, making them lightweight add-ons. In text‐only LLMs, such methods aim to align outputs with factual knowledge [chuang2023dola, li2023inference]. In MLLMs, post-decoding strategies emphasize the role of visual inputs [leng2024mitigating, zhao2024mitigating, favero2024multi] and can be classified into image-level and token-level interventions. Image-level decoding methods [zhang2024debiasing, manevich2024mitigatinghallucinationslargevisionlanguage] treat all objects in the input image uniformly, limiting their effectiveness in addressing co-occurrence hallucinations. Existing token-level methods either rely on external segmentation [chen2024halc] and detection models [kan2024catchcomplementaryadaptivetokenlevel] or lack awareness of object-related decoupling [woo2024don]. Moreover, these methods typically introduce an implicit trade-off between accuracy and informativeness, reducing hallucinations at the expense of omitting valid details. Attention-based methods [tang2025intervening, zhang2024redundancy] require careful selection of specific layers and often introduce model-specific adjustments or heuristics. In contrast, our GACD directly estimates embedded bias and decouples object-aware visual tokens, enabling sample-specific hallucination mitigation without external data, models, or model-specific adjustments, while achieving a more favorable balance between accuracy and informativeness.
@@ A MLLM with parameters $\theta$ computes, at each decoding step $m$, a logit vector
Given a dataset $\mathcal{D}$ of $(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y})$, maximum-likelihood training (or fine-tuning) estimates $\theta^{\star}$ by maximizing the conditional log-likelihood. Pretrained MLLMs encode statistical regularities (including spurious correlations) from training data in $\theta^{\star}$; such behavior can be probed without changing $\theta^{\star}$ via parameter-dependent analyses (e.g., gradients/attributions or counterfactual decodings) [kim2019learning, guo2024bias], enabling self-reflective bias interpretation.
@@ 3.2 Gradient-Based Token Influence Estimation
To capture these embeded biases, we measure how each input token perturbs the output logits. Let $\mathbf{z}^{\star}_{m}\in\mathbb{R}^{|\mathcal{V}|}$ denote the step-$m$ logits $\mathbf{z}^{\star}_{m}=\pi_{\theta^{\star}}(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m})$. Around a reference sample point $(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}_{<m}^{(0)})$, the first-order Taylor expansion [spivak1980calculus] of the logits $\mathbf{z}^{\star}_{m}$ is
@@
where $\big|_{\cdot}$ indicate evaluation at the reference sample point. Taylor expansion details are in supplementary Sec. 1. Each $\mathbf{g}^{v}_{ms},\,\mathbf{g}^{p}_{mn},\,\mathbf{g}^{y}_{mi}$ indicate a small token perturbation in its corresponding embedding space to a perturbation of the predict logit vector in $\mathbb{R}^{|\mathcal{V}|}$. We score the importance of each input token by the Manhattan norm of its gradient:
@@ 3.3 Influence-Aware Constrained Decoding
GACD builds on token influence estimation with two components: (i) Object-aware Visual Token Grouping and (ii) Anchor-specific Influence-weighted Decoding. At step $m$, the former partitions visual tokens into object-related $\mathbf{t}^{o}$ and unrelated $\mathbf{t}^{u}$ based on objects detected in $\mathbf{y}_{<m}$. The latter extends contrastive decoding [li2022contrastive] by forming Anchor-specific negative guidance logits from pre-mentioned objects and computing a decoding weight $\alpha_{m}$ from token-influence measurements.
@@
Unlike existing decoding methods [zhou2023analyzing, favero2024multi, leng2024mitigating], which rely on adaptive plausibility constraints (e.g., prediction confidence) and require experimental tuning to determine optimal thresholds, our approach explicitly enforces non-negativity on the influence of object-related visual and prompt tokens. This corresponds to the following upper-bound condition:
@@
Additionally, since hallucinations are more likely in long generations [zhou2023analyzing, peng2025mitigating], we introduce a sample-dependent stopping criterion based on visual influence. Specifically, if the visual influence ratio $r^{v}_{m}$ of the token following the end-of-sequence (EOS) falls below a threshold $\epsilon$,
@@ 4 Experiments
The proposed method is evaluated for both the open-ended generative tasks and the discriminative tasks. We use Amber [wang2023llm], MSCOCO [lin2014microsoft] and LLaVa-QA90 [liu2024visual] datasets for the generative task, and Amber [wang2023llm] and POPE [li2023evaluating] datasets on the discriminative tasks.
@@ 4.2 Implementation Details
The maximum output length is set to 256 across all models, with other model parameters kept at their defaults. To prevent excessive modifications, we set the maximum amplification factor, $\alpha_{m}$, to 5 for discriminative tasks and 3 for generative tasks. We empirically set the early stopping thresholds $\epsilon$ as follows: LLaVA-v1.5 and LLaVA-v1.6: $7\%$, InstrucBLIP: $25\%$, mPLUG-Owl2: $2.5\%$, and InternVL2: $10\%$. All experiments are performed on an NVIDIA A40 GPU with batch size of 1. Unless otherwise specified, we use greedy sampling [graves2013generating].
@@ 4.3 Results on Open-ended Generation
RLAIFv yu2024rlaif 6.6 49.7 32.0 2.9 76.7 92.0 78.1 84.5 89.0
VCD leng2024mitigating 6.7 46.4 32.6 3.5 71.3 91.1 62.3 74.3 83.8
M3ID favero2024multi 6.2 50.5 29.3 2.8 72.4 91.8 64.1 75.5 84.7
AVISC woo2024don 6.5 50.2 34.8 2.7 73.8 89.7 68.4 77.6 85.5
@@
RLAIFv yu2024rlaif 7.6 47.7 29.9 2.8 76.5 84.5 79.0 81.7 87.1
VCD leng2024mitigating 7.9 49.7 36.7 3.7 75.9 83.5 79.3 81.3 86.7
M3ID favero2024multi 7.3 49.2 33.8 3.7 75.8 84.4 77.9 81.0 86.9
AVISC woo2024don 7.1 48.8 34.4 4.3 75.9 83.4 79.5 81.4 87.2
@@
RLAIFv yu2024rlaif 7.8 50.5 35.7 4.1 81.2 90.8 79.7 84.9 88.6
VCD leng2024mitigating 8.0 51.3 35.3 4.1 75.6 83.5 78.8 81.1 86.6
M3ID favero2024multi 7.8 51.7 34.9 4.1 75.9 83.5 79.3 81.3 86.8
AVISC woo2024don 10.9 50.5 35.5 4.4 82.1 90.7 81.4 85.8 87.5
@@
Table 2: Open-ended Generation Results Using the CHAIR Metric on the MSCOCO Subset Following [deng2024seeing].
@@
Table 3: Results on LLaVA-QA90 Dataset, with All Settings Following [leng2024mitigating].
@@ 4.4 Results on Discriminative Task
Variation in Improvement Across Question Categories. Fig. 3(b) presents F1 scores across various question categories using LLaVA-v1.5 [liu2024llava]. Our method improves performance across all categories, with the largest gains in existence, attributes, and state—categories strongly tied to visual cues, benefiting from enhanced visual token influence.
@@ 4.5 Results on Additional MLLMs and Datasets
We further evaluate modern MLLMs, LLaVA-v1.6, Qwen2-VL and Intern-VL2, on the Amber dataset. As shown in Tab. 1 in supplementary, our method consistently improves performance, even for strong baselines. To further evaluate the effectiveness of our method across diverse tasks, we conduct evaluations on the MMBench [liu2025mmbench], MMVet [yu2023mm], ScienceQA [lu2022learn] and VizWiz [gurari2018vizwiz] datasets. Results in supplementary show that our approach enhances perception and recognition, especially for tasks with clear visual grounding, while preserving overall task performance and demonstrating robustness.
@@ 5.2 Co-occurrence Bias Analysis
Fig. 5(a) shows an example where mPLUG-Owl2 incorrectly predicts ‘dining table’ due to the presence of a ‘chair’. In Fig. 5(b), the influence of individual visual tokens on hallucinated prediction $I_{ms}(\text{`table'})$ and $I_{ms}(\text{`chair'})$ shows that they share the same most influential visual token: $s=24$. These visualizations indicate that the influence distribution over tokens is typically sharply peaked, so selecting the single most influential token gives a clean and interpretable attribution signal while reducing noise from low-influence tokens. We further collected 100 chair-only and 100 table-only images from MSCOCO evaluation dataset [lin2014microsoft].

### v4

@@ 1 Introduction
Recent advances in Multimodal Large Language Models (MLLMs) show strong ability to produce coherent and context-aware content across a wide range of domains [2, 9, 6, 31, 51]. Despite their impressive advancements, these models remain prone to hallucination, wherein the generated text is not faithfully grounded in the visual modality [42, 29]. This limitation poses a critical barrier to establishing trust in the outputs of MLLMs.
@@
The hallucinations observed in MLLMs can be largely attributed to two fundamental biases [22, 29, 25]. Text-visual bias refers to the excessive reliance on textual information—such as the input prompt and previously generated outputs—while neglecting the visual modality during generation. This bias becomes particularly pronounced in longer sequences, where MLLMs tend to depend more heavily on prior text and increasingly disregard visual cues [61, 12]. Co-occurrence bias arises from spurious statistical correlations embedded in the training data, which lead models to erroneously predict the presence of non-existent objects based on their frequent co-occurrence with observed objects in the visual inputs [29]. This bias is particularly challenging to mitigate, and existing approaches largely rely on statistical priors rather than offering statistically agnostic solutions [22, 60].
@@
Existing efforts to mitigate hallucinations in MLLMs can be broadly categorized into inference-based methods, which operate at the decoding stage [5, 12, 26, 39, 49], and training-based approaches, which intervene during model optimization [3, 7, 22, 45, 19, 54]. Inference-based approaches are valued for their cost-effectiveness, as they avoid the need for additional data collection, data bias examination, or extensive model retraining. However, these methods offer limited insight into the severity of underlying biases, leaving the root causes of hallucination insufficiently understood. In addition, some inference-based methods rely on auxiliary models—such as segmentation networks  [5], detection systems  [21] , or even additional MLLMs [41, 10, 50]— which introduce extra sources of error, depend on task-specific supervision.
@@
Another limitation of existing methods lies in their lack of granularity when adjusting the underlying biases in MLLMs. Most approaches rely on heuristically tuned priors, which vary across datasets and fail to generalize reliably [26, 59]. Moreover, they apply uniform weighting across all visual features, offering no mechanism to selectively adjust bias at the level of individual features [58, 37]. This coarse treatment limits their effectiveness in mitigating co-occurrence bias, which arises from spurious statistical correlations between objects that are often represented by distinct visual features.
@@
  • Extensive experiments demonstrate that GACD mitigates hallucinations and enhances accuracy while maintaining a favorable balance between accuracy and informativeness. GACD achieves up to $8\%$ increase in overall score on AMBER [48], an $8\%$ F1 boost on POPE [29], up to $45\%$ improvement in detailness and a $92\%$ accuracy gain on LLaVA-QA90 [32].
@@ 2 Related Work
Hallucination and Bias. Hallucinations in LLMs often arise from biases in the training data [38, 17], while in MLLMs, studies [47, 29, 13] show that hallucinations are closely linked to biases like text-visual and co-occurrence biases. Additionally, biases related to output position, which increase the risk of hallucination as output length grows, have been examined in [12, 61]. Existing methods [29, 13, 25] typically report only overall statistics, lacking a mathematical, sample-wise bias measurement. This distinction is important, as biases can vary case by case. Our approach measures sample-dependent bias via token-level gradient sensitivities, revealing how pre-trained MLLM parameters embed these biases  [24, 15], and enabling self-reflective hallucination mitigation.
Hallucination Mitigation. Training-related hallucination mitigation methods [4, 19, 54, 40, 55] are expensive, requiring access to training data and specialized statistical analysis. Among them, LPOI [55] also employs an object-aware framework, highlighting the effectiveness of modeling object-level information for mitigating hallucinations. Reinforcement‐learning approaches [50, 10, 56] rely on supplementary feedback, often from human annotators or auxiliary LLMs/MLLMs, and the latter may themselves hallucinate. By contrast, post-decoding techniques modify model logits at inference time without further training or external feedback, making them lightweight add-ons. In text‐only LLMs, such methods aim to align outputs with factual knowledge [8, 27]. In MLLMs, post-decoding strategies emphasize the role of visual inputs [26, 59, 12] and can be classified into image-level and token-level interventions. Image-level decoding methods [58, 37] treat all objects in the input image uniformly, limiting their effectiveness in addressing co-occurrence hallucinations. Existing token-level methods either rely on external segmentation [5] and detection models [21] or lack awareness of object-related decoupling [49]. Moreover, these methods typically introduce an implicit trade-off between accuracy and informativeness, reducing hallucinations at the expense of omitting valid details. Attention-based methods [46, 57] require careful selection of specific layers and often introduce model-specific adjustments or heuristics. In contrast, our GACD directly estimates embedded bias and decouples object-aware visual tokens, enabling sample-specific hallucination mitigation without external data, models, or model-specific adjustments, while achieving a more favorable balance between accuracy and informativeness.
@@ A MLLM with parameters $\theta$ computes, at each decoding step $m$, a logit vector
Given a dataset $\mathcal{D}$ of $(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y})$, maximum-likelihood training (or fine-tuning) estimates $\theta^{\star}$ by maximizing the conditional log-likelihood. Pretrained MLLMs encode statistical regularities (including spurious correlations) from training data in $\theta^{\star}$; such behavior can be probed without changing $\theta^{\star}$ via parameter-dependent analyses (e.g., gradients/attributions or counterfactual decodings) [24, 15], enabling self-reflective bias interpretation.
@@ 3.2 Gradient-Based Token Influence Estimation
To capture these embeded biases, we interpret how each input token perturbs the output logits. Let $\mathbf{z}^{\star}_{m}\in\mathbb{R}^{|\mathcal{V}|}$ denote the step-$m$ logits $\mathbf{z}^{\star}_{m}=\pi_{\theta^{\star}}(\mathbf{t}^{v},\mathbf{t}^{p},\mathbf{y}_{<m})$. Around a reference sample point $(\mathbf{t}^{v(0)},\mathbf{t}^{p(0)},\mathbf{y}_{<m}^{(0)})$, the first-order Taylor expansion [44] of the logits $\mathbf{z}^{\star}_{m}$ is
@@
where $\big|_{\cdot}$ indicate evaluation at the reference sample point. Taylor expansion details are in supplementary Sec. 1. Each $\mathbf{g}^{v}_{ms},\,\mathbf{g}^{p}_{mn},\,\mathbf{g}^{y}_{mi}$ indicate a small token perturbation in its embedding space to a perturbation of the predict logit vector in $\mathbb{R}^{|\mathcal{V}|}$. Following [43], we approximate the importance of each input token by the Manhattan norm of its gradient:
@@ 3.3 Influence-Aware Constrained Decoding
GACD builds on token influence estimation with two components: (i) Object-aware Visual Token Grouping and (ii) Anchor-specific Influence-weighted Decoding. At step $m$, the former partitions visual tokens into object-related $\mathbf{t}^{o}$ and unrelated $\mathbf{t}^{u}$ based on objects detected in $\mathbf{y}_{<m}$. The latter extends contrastive decoding [28] by forming Anchor-specific negative guidance logits from pre-mentioned objects and computing a decoding weight $\alpha_{m}$ from token-influence measurements.
@@
Unlike existing decoding methods [61, 12, 26], which rely on adaptive plausibility constraints (e.g., prediction confidence) and require experimental tuning to determine optimal thresholds, our approach explicitly enforces non-negativity on the influence of object-related visual and prompt tokens. This corresponds to the following upper-bound condition:
@@
Additionally, since hallucinations are more likely in long generations [61, 40], we introduce a sample-dependent stopping criterion based on visual influence. Specifically, if the visual influence ratio $r^{v}_{m}$ of the token following the end-of-sequence (EOS) falls below a threshold $\epsilon$,
@@ 4 Experiments
The proposed method is evaluated for both the open-ended generative tasks and the discriminative tasks. We use Amber [48], MSCOCO [30] and LLaVa-QA90 [32] datasets for the generative task, and Amber [48] and POPE [29] datasets on the discriminative tasks.
@@ 4.2 Implementation Details
The maximum output length is set to 256 across all models, with other model parameters kept at their defaults. Gradients are obtained via PyTorch’s torch.autograd.grad on the input tokens. Noun tokens are identified using the spaCy library via its en_core_web_sm model. To prevent excessive modifications, we set the maximum amplification factor, $\alpha_{m}$, to 5 for discriminative tasks and 3 for generative tasks. We empirically set the early stopping thresholds $\epsilon$ as follows: LLaVA-v1.5 and LLaVA-v1.6: $7\%$, InstrucBLIP: $25\%$, mPLUG-Owl2: $2.5\%$, and InternVL2: $10\%$. All experiments are performed on an NVIDIA A40 GPU with batch size of 1. Unless otherwise specified, we use greedy sampling [14].
@@ 4.3 Results on Open-ended Generation
RLAIFv 52 6.6 49.7 32.0 2.9 76.7 92.0 78.1 84.5 89.0
VCD 26 6.7 46.4 32.6 3.5 71.3 91.1 62.3 74.3 83.8
M3ID 12 6.2 50.5 29.3 2.8 72.4 91.8 64.1 75.5 84.7
AVISC 49 6.5 50.2 34.8 2.7 73.8 89.7 68.4 77.6 85.5
@@
RLAIFv 52 7.6 47.7 29.9 2.8 76.5 84.5 79.0 81.7 87.1
VCD 26 7.9 49.7 36.7 3.7 75.9 83.5 79.3 81.3 86.7
M3ID 12 7.3 49.2 33.8 3.7 75.8 84.4 77.9 81.0 86.9
AVISC 49 7.1 48.8 34.4 4.3 75.9 83.4 79.5 81.4 87.2
@@
RLAIFv 52 7.8 50.5 35.7 4.1 81.2 90.8 79.7 84.9 88.6
VCD 26 8.0 51.3 35.3 4.1 75.6 83.5 78.8 81.1 86.6
M3ID 12 7.8 51.7 34.9 4.1 75.9 83.5 79.3 81.3 86.8
AVISC 49 10.9 50.5 35.5 4.4 82.1 90.7 81.4 85.8 87.5
@@
Table 2: Open-ended Generation Results Using the CHAIR Metric on the MSCOCO Subset Following [10].
@@
Table 3: Results on LLaVA-QA90 Dataset, with All Settings Following [26].
@@ 4.4 Results on Discriminative Task
Variation in Improvement Across Question Categories. Fig. 3(b) presents F1 scores across various question categories using LLaVA-v1.5 [31]. Our method improves performance across all categories, with the largest gains in existence, attributes, and state—categories strongly tied to visual cues, benefiting from enhanced visual token influence.
@@ 4.5 Results on Additional MLLMs and Datasets
We further evaluate modern MLLMs, LLaVA-v1.6, Qwen2-VL and Intern-VL2, on the Amber dataset. As shown in Tab. 1 in supplementary, our method consistently improves performance, even for strong baselines. To further evaluate the effectiveness of our method across diverse tasks, we conduct evaluations on the MMBench [34], MMVet [53], ScienceQA [35] and VizWiz [16] datasets. Results in supplementary show that our approach enhances perception and recognition, especially for tasks with clear visual grounding, while preserving overall task performance and demonstrating robustness.
@@ 5.2 Co-occurrence Bias Analysis
Fig. 5(a) shows an example where mPLUG-Owl2 incorrectly predicts ‘dining table’ due to the presence of a ‘chair’. In Fig. 5(b), the influence of individual visual tokens on hallucinated prediction $I_{ms}(\text{`table'})$ and $I_{ms}(\text{`chair'})$ shows that they share the same most influential visual token: $s=24$. These visualizations indicate that the influence distribution over tokens is typically sharply peaked, so selecting the single most influential token gives a clean and interpretable attribution signal while reducing noise from low-influence tokens. We further collected 100 chair-only and 100 table-only images from MSCOCO evaluation dataset [30].